OpenStudy (anonymous):

Please explain! (: If we toss 3 coins 100,000 times. How many times do we expect to see "Exactly 1 Head" appearing? A.12,500 B.37,500 C.3,750 D.50,000

OpenStudy (anonymous):

12500

OpenStudy (anonymous):

can you explain why?

OpenStudy (anonymous):

is it not 37500? P( head) = 0.5 HTT THT TTH 3 situations, with same probability: 3*(0.5*0.5*0.5) = 3/8 Expected = 3/8*100000 = 37500

OpenStudy (anonymous):

(3/8)*100000

OpenStudy (anonymous):

the probability of a head on one toss = 1/2 if we have one head on tossing 3 coins together possibilities are htt tht tth probabilty of any of these = 3 * 1/8 = 3/8 so the answer is 3/8 * 100,000 = 37500

OpenStudy (hoblos):

for each toss we have 8 cases (hhh)(hht)(hth)(thh)(htt)(tht)(tth)(ttt) 3 of these have only 1 head so the probability of obtaining only 1 head is 3/8 for 100,000 tosses we expect to see 100000(3/8) = 37500